# Classifying representations of finite classical groups of Lie type of dimension up to $\ell^{4}$

### v1

@@ 4. Representations of finite groups of type $A_{\ell}$
Liberty of coefficients -- Assuming that $a_{\ell-3}\neq 0$ we shall show that $a_{\ell-3}=1$ and $a_{i}=0$ for any $i\leq 4$ and $i\geq\ell-2$. In fact, if $a_{\ell-j}\neq 0$ for all $j=0,1,2$ then the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-4}$. We then see that
(4.1) $\dim V(\lambda)\geq l(\lambda)=\frac{(\ell+1)!}{(\ell-3)!}=(\ell-2)(\ell-1)\ell(\ell+1)>\frac{\ell^{4}}{16}.$
On the other hand, when exactly two coefficients $a_{\ell-j}$ are nonzero, the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-3}\times A_{2}$. An argument as above shows that $\dim V(\lambda)>\ell^{4}/16$. Next, if just one coefficient $a_{\ell-j}$ is nonzero then the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-4}\times A_{1}\times A_{1}$ or $A_{\ell-4}\times A_{2}$ according to $j=1$ or not. Thus, we may rule out this possibility of $\lambda$. It follows that $a_{\ell-j}=0$ for all $j=0,1,2$ and by a symmetry argument one checks that $a_{i}=0$ for all $i=1,2,3$. Thus, we reduce to $\lambda=a_{\ell-3}\lambda_{\ell-3}$.
We now will prove that $a_{\ell-3}=1$ by supposing the contrary i.e., $a_{\ell-3}\geq 2$. Then, we observe that the $(\ell-4)$-coefficient of
@@
We shall prove that $a_{i}=0$ for $1\leq i\leq 4$. Suppose first $a_{1}\neq 0$ and $a_{2}=a_{3}=a_{4}=0$. We note that the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-4}$ or $A_{\ell-4}\times A_{1}$ or $A_{\ell-4}\times A_{2}$ according to $a_{\ell-1}\neq 0\neq a_{\ell}$ or not. Then, reasoning as above, we can deduce that $\dim V(\lambda)>\ell^{4}/16$. Similarly each possibility of some $a_{i}\neq 0$ for $1\leq i\leq 4$ discards $\lambda$ as admissible weight, and thus
@@
We now show that $a_{\ell-2}\leq 1$. Otherwise, assuming that $a_{\ell-2}\geq 2$ one observes that the coefficients at $(\ell-3)$ and $(\ell-1)$ of the dominant weight $\lambda-a_{\ell-2}$ are nonzero, but this is not possible by reduction above. Hence, we reduce our analysis to study
$\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}\quad\text{or }\quad\lambda=a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
When $a_{\ell-2}=1$ and $a_{\ell-1}\geq 2$ then the term $2\lambda_{\ell-2}$ appears in $\lambda-\alpha_{\ell-1}$, which is impossible by applying the reasoning above. So, any $\lambda$ of the form $\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ with $a_{\ell-1}\geq 2$ is not admissible and we reduce us to $\lambda$ of the form $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ or $\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}$. For the case $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$, if $a_{\ell}\geq 2$ we may observe similarly that the term $2\lambda_{\ell-1}$ appears in the dominant weight $\lambda-\alpha_{\ell-1}$ which is a contradiction by iterating the previous reasoning. This reduces our analysis to study the dominant weights $\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ and $\lambda_{\ell-2}+\lambda_{\ell-1}$. In the same vein, we discard the possibility of a dominant weight of the form $\lambda_{\ell-2}+a_{\ell}\lambda_{\ell}$ with $a_{\ell}\geq 4$. The cases $a_{\ell}=2,3$ are reduced again to analyze the dominant weights $\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ and $\lambda_{\ell-2}+\lambda_{\ell-1}$. Finally, when $\lambda=\lambda_{\ell-2}+a_{\ell}\lambda_{\ell}$ with $a_{\ell}\geq 2$ cannot be admissible by using similar arguments as above. Therefore, we only need to determinate the dimensions of $V(\lambda)$ for the following dominants weights
@@
$\dim V(\lambda_{\ell-2}+\lambda_{\ell-1})\geq l(\lambda_{\ell-3}+\lambda_{\ell})$ $=\frac{\ell^{4}(1-2/\ell)(1-1/\ell)(1+1/\ell)}{6}>\frac{\ell^{4}(1-2/\ell)(1-1/\ell)}{6}>\ell^{4}/12>\ell^{4}/16,$
since $(1-2/\ell)(1-1/\ell)\geq((1-2/15)(1-1/15)>1/2$ for any $\ell\geq 15$. In the same way, we can rule out the second case by using that $\lambda_{\ell-3}+2\lambda_{\ell}\prec\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$. Note that $\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible by appealing to Example 3.4 and $\lambda_{\ell-2}$ is admissible by Proposition 3.6.
@@
If $a_{\ell-1}\geq 2$ we note again that $\lambda_{\ell-2}$ and $\lambda_{\ell}$ appear in the writing of the dominant weight $\lambda-\alpha_{\ell-1}$, which is impossible as was showed previously. Thus, we reduce to $a_{\ell-1}=1$, i.e.,
@@
In this case we may prove that $a_{i}=0$ for $2\leq i\leq 4$. Here, many cases arise for the $\mathcal{W}$-stabilizer of $\lambda$ according to the vanished for some of these $a_{i}$ ($2\leq i\leq 4)$. For instance, if $a_{2}\neq 0$ and $a_{3}=a_{4}=0$ we see that $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{1}\times A_{\ell-4}$ or $A_{\ell-4}$ according to $a_{1}=0$ or not. Thus, we may verify that $\dim V(\lambda)>\ell^{4}/16$, i.e., $\lambda$ is not admissible. Indeed, in any case we may deduce that the dimension $V(\lambda)$ is bigger than the bound $\ell^{4}/16$. This implies
@@
Reasoning as above one verifies that $a_{1}$ and $a_{\ell}$ cannot be simultaneously nonzero. In fact, assume first that $a_{1}\geq 2$ we obtain the dominant weight
@@
which is ruled out by computations as above. On the other hand, if $a_{\ell}\geq 2$ then $2\lambda_{\ell-1}$ appears in the writing of $\lambda-\alpha_{\ell}$, which is impossible too. Moreover, we may discard $\lambda=\lambda_{1}+\lambda_{\ell-1}+\lambda_{\ell}$ by arguing as in Example 3.4. This lefts us the dominant weights $\lambda_{1}+\lambda_{\ell-1}$, $\lambda_{\ell-1}+\lambda_{\ell}$ and $\lambda_{\ell-1}$ which are admissible by Proposition 3.6.
Finally suppose that $a_{\ell-3}=a_{\ell-2}=a_{\ell-1}=0$ and $a_{\ell}\neq 0$, i.e., we will deal with
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+a_{3}\lambda_{3}+a_{4}\lambda_{4}+a_{\ell}\lambda_{\ell}$
@@
Observe that if $a_{2}\neq 0$ then $a_{1}=0$ by using a dual reasoning to the previous case. This implies that $\lambda=\lambda_{2}+a_{\ell}\lambda_{\ell}$, which gives just one new admissible weight $\lambda=\lambda_{2}+\lambda_{\ell}$. Finally, if $a_{2}=0$ then we have that $\lambda=a_{1}\lambda_{1}+a_{\ell}\lambda_{\ell}$ and we get the new admissible weights $\lambda_{1}+\lambda_{\ell}$, $2\lambda_{1}+\lambda_{\ell}$, $\lambda_{1}+2\lambda_{\ell}$ and $a_{\ell}\lambda_{\ell}$ for $a_{\ell}=1,2,3,4$.
In summary, we have proved the first of our main results:
Theorem 4.1.
Let $G$ be a finite group of Lie type $A_{\ell}$ over $\mathbb{\overline{F}}_{p}$. If $\ell\geq 15$ then the list giving in Proposition 3.6 exhausts the admissible weights.
@@ 5. Multiplicities and dimension in finite classical groups
In this section we deal with the rest of finite classical groups. Let $G$ be a connected reductive simple algebraic group of type $B_{\ell}$, $C_{\ell}$ or $D_{\ell}$ of simply connected type. By analogy to Definition 3.2 we present the notion of admissibility for these types.
Definition 5.1.
Let $G$ be a connected reductive simple algebraic group of type $B_{\ell}$, $C_{\ell}$ or $D_{\ell}$ over $\overline{\mathbb{F}}_{p}$ of simply connected type and $\lambda$ be a dominant weight. We will say that $\lambda$ is admissible if the dimension of the $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ is at most $\ell^{4}$.
We shall begin the desired classification of admissible weights by considering $G$ of type $B_{\ell}$.
@@ 5.1. Type $B_{\ell}$
In order to compute the multiplicities involved in this section, we start by considering $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{\ell}\}$ the simple roots system (of type $B_{\ell}$) given by
@@
(5.1) $(\alpha_{i},\alpha_{j})=\begin{cases}\phantom{-}2,&\text{ if }i=j\\ -1,&\text{ if }i=j\pm 1\\ \phantom{-}0,&\text{ otherwise }.\\ \end{cases}$
@@
$(\alpha_{i},\delta)=\begin{cases}\frac{1}{2},&\text{ if }i=1\\ 1,&\text{ otherwise }.\\ \end{cases}$
@@
For all $\ell\geq 12$, the weight $\lambda=4\lambda_{\ell}$ is admissible. To see that, it suffices to show that $\dim\operatorname{Sym}^{4}(V)$ is less or equal to $\ell^{4}$ since $V(4\lambda_{\ell})\subseteq\operatorname{Sym}^{4}(V)$. In fact, we observe that
$\dim\operatorname{Sym}^{4}(V)$ $=\frac{(2\ell+1)(2\ell+2)(2\ell+3)(2\ell+4)}{24}$
$=\frac{(2\ell+1)(\ell+1)(2\ell+3)(\ell+2)}{6}$
$=\ell^{4}\cdot\frac{(2+1/\ell)(1+1/\ell)(2+3/\ell)(1+2/\ell)}{6}$
$\leq\ell^{4},$
@@
To find all the admissible weights $\lambda$ and compute dimensions of certain $G$-modules $V(\lambda)$, we will need the following lemmas.
@@
For all $\ell\geq 12$, the multiplicities of all dominant weights $\mu\preceq i\lambda_{\ell}$, $i=3,4$, are given in the following table:
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$
$3\lambda_{\ell}$ $3\lambda_{\ell}$ $1$
$\lambda_{\ell-1}+\lambda_{\ell}$ $1$
$\lambda_{\ell-2}$ $1$
$2\lambda_{\ell}$ $1$
$\lambda_{\ell-1}$ $1$
[…]
00 $\ell$
$4\lambda_{\ell}$ $4\lambda_{\ell}$ $1$
$\lambda_{\ell-1}+2\lambda_{\ell}$ $1$
$2\lambda_{\ell-1}$ $1$
$\lambda_{\ell-2}+\lambda_{\ell}$ $1$
$\lambda_{\ell-3}$ $1$
$3\lambda_{\ell}$ $1$
$\lambda_{\ell-1}+\lambda_{\ell}$ $1$
$\lambda_{\ell-2}$ $1$
[…]
00 $\ell(\ell-1)/2$
Proof.
Let $\lambda=3\lambda_{\ell}$. We shall compute those multiplicities $m_{\lambda}(\mu)$ where $2\lambda_{\ell}\prec\mu\preceq 3\lambda_{\ell}$. Since $\mu:=\lambda_{\ell-1}+\lambda_{\ell}=\lambda-\alpha_{\ell}$ we have that $(\lambda,\lambda)=9$ and $(\mu,\mu)=5$. In addition as $(\alpha_{\ell},\delta)=1$ and $(\lambda,\alpha_{\ell})=3(\lambda_{\ell},\alpha_{\ell})=1$, we have
Next, we compute the multiplicity of $\lambda_{\ell-2}=\lambda-(\alpha_{\ell-1}+2\alpha_{\ell})$. To do this, we observe that the positive roots involved in Fruedenthal formula are $\alpha_{\ell-1}$, $\alpha_{\ell}$, $2\alpha_{\ell}$ and $\alpha_{\ell-1}+\alpha_{\ell}$. Notice, at this point, that $\lambda$ is not included in the formula since $\mu+\alpha_{\ell-1}+2\alpha_{\ell}$ is not of the form $\mu+i\alpha$ for some positive root $\alpha$ by appealing to [3, Plate II].
Set $\mu:=\lambda-(\alpha_{\ell-1}+\alpha_{\ell})$.
  1. (1)
    $m_{\lambda}(\mu_{\ell-1})=2\cdot\frac{m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell})(\mu_{\ell-1}+\alpha_{\ell},\alpha_{\ell})}{(\lambda,\lambda)-(\mu_{\ell-1},\mu_{\ell-1})+2(\alpha_{\ell},\delta)}=2\cdot\frac{1\cdot 3}{9-5+2}=1.$
  2. (2)
$m_{\lambda}(\mu)$ $=2\cdot\frac{m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell-1})(\mu_{\ell-1}+\alpha_{\ell-1},\alpha_{\ell-1})+m_{\lambda}(\mu_{\ell}+\alpha_{\ell})(\mu_{\ell}+\alpha_{\ell},\alpha_{\ell})+m_{\lambda}(\lambda)(\lambda_{\ell},\alpha_{\ell-1}+\alpha_{\ell})}{(\lambda,\lambda)-(\lambda_{\ell-2},\lambda_{\ell-2})+2(\alpha_{\ell-1}+\alpha_{\ell},\delta)}=1$
$=2\cdot\frac{1\cdot 1+0+1\cdot 3}{9-5+4}=1.$
$m_{\lambda}(\lambda-2\alpha_{\ell})$ $=2\cdot\frac{m_{\lambda}(\lambda-\alpha_{\ell})(\lambda-\alpha_{\ell},\alpha_{\ell})+m_{\lambda}(\lambda)(\lambda,\alpha_{\ell})}{(\lambda,\lambda)-(\mu,\mu)+2(2\alpha_{\ell},\delta)}$
$=2\cdot\frac{1\cdot 1+1\cdot 3}{9-5+4}=1.$
where
(*(*)( * ) $=(\mu+\alpha_{\ell},\alpha_{\ell})m_{\lambda}(\mu+\alpha_{\ell})+(\mu+\alpha_{\ell-1})m_{\lambda}(\mu+\alpha_{\ell-1},\alpha_{\ell-1})$
Computing as above we see that
$m_{\lambda}(\mu)=\frac{2[1\cdot 2+1\cdot 2+0\cdot 4+1\cdot 2]}{12}=1.$
The remaining multiplicities $m_{\lambda}(\mu)$, with $\mu\preceq 2\lambda$ were compute in [11].
Finally, the case $\lambda=4\lambda_{\ell}$ is longer but completely similar than the one above by computing the multiplicity of $\lambda_{\ell-1}+2\lambda_{\ell}$ as first step. ∎
Lemma 5.4.
For all $\ell\geq 12$, the multiplicities of all dominant weights $\mu\preceq\lambda_{\ell-i}$, $i=2,3$, are given in the following table:
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$
$\lambda_{\ell-2}$ $\lambda_{\ell-2}$ $1$
$\lambda_{\ell-1}$ $1$

### v3

@@ 3. Representations of finite classical groups of type $A_{\ell}$
Liberty of coefficients - Assuming that $a_{\ell-3}\neq 0$ we shall show that $a_{\ell-3}=1$ and $a_{i}=0$ for any $i\leq 4$ and $i\geq\ell-2$. In fact, if $a_{\ell-j}\neq 0$ for all $j=0,1,2$ then the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-4}$. We then see that
(3.2) $\dim L(\lambda)\geq l(\lambda)=\frac{(\ell+1)!}{(\ell-3)!}=(\ell-2)(\ell-1)\ell(\ell+1)>\frac{\ell^{4}}{16}.$
On the other hand, when exactly two coefficients $a_{\ell-j}$ are nonzero, the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-3}\times A_{2}$. An argument as above shows that $\dim L(\lambda)>\ell^{4}/16$. Next, if just one coefficient $a_{\ell-j}$ is nonzero then the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-4}\times A_{1}\times A_{1}$ or $A_{\ell-4}\times A_{2}$ according to $j=1$ or not. Thus, we may rule out this possibility of $\lambda$. It follows that $a_{\ell-j}=0$ for all $j=0,1,2$ and by a symmetry argument one checks that $a_{i}=0$ for all $i=1,2,3$. Thus, we reduce to $\lambda=a_{\ell-3}\lambda_{\ell-3}$. We will now prove that $a_{\ell-3}=1$ by supposing the contrary i.e., $a_{\ell-3}\geq 2$. Then, we observe that the $(\ell-4)$-coefficient of
@@
We shall prove that $a_{i}=0$ for $1\leq i\leq 4$. Suppose first $a_{1}\neq 0$ and $a_{2}=a_{3}=a_{4}=0$. We note that the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-4}$ or $A_{\ell-4}\times A_{1}$ or $A_{\ell-4}\times A_{2}$ according to $a_{\ell-1}\neq 0\neq a_{\ell}$ or not. Then, reasoning as above, we can deduce that $\dim L(\lambda)>\ell^{4}/16$. Similarly each possibility of some $a_{i}\neq 0$ for $1\leq i\leq 4$ discards $\lambda$ as an admissible weight, and thus
@@
We now show that $a_{\ell-2}\leq 1$. Otherwise, assuming that $a_{\ell-2}\geq 2$, one observes that the coefficients at $(\ell-3)$ and $(\ell-1)$ of the dominant weight $\lambda-a_{\ell-2}$ are nonzero, but this is not possible by reduction above. Hence, we reduce our analysis to study
[…]
When $a_{\ell-2}=1$ and $a_{\ell-1}\geq 2$ then the term $2\lambda_{\ell-2}$ appears in $\lambda-\alpha_{\ell-1}$, which is impossible by applying the reasoning above. So, any $\lambda$ of the form $\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ with $a_{\ell-1}\geq 2$ is not admissible and we reduce us to $\lambda$ of the form $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ or $\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}$. For the case $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$, if $a_{\ell}\geq 2$ we may observe similarly that the term $2\lambda_{\ell-1}$ appears in the dominant weight $\lambda-\alpha_{\ell-1}$ which is a contradiction by iterating the previous reasoning. This reduces our analysis to study the dominant weights $\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ and $\lambda_{\ell-2}+\lambda_{\ell-1}$. In the same vein, we discard the possibility of a dominant weight of the form $\lambda_{\ell-2}+a_{\ell}\lambda_{\ell}$ with $a_{\ell}\geq 4$. The cases $a_{\ell}=2,3$ are reduced again to analyze the dominant weights $\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ and $\lambda_{\ell-2}+\lambda_{\ell-1}$. Finally, when $\lambda=\lambda_{\ell-2}+a_{\ell}\lambda_{\ell}$ with $a_{\ell}\geq 2$ cannot be admissible by using similar arguments as above. Therefore, we only need to determine the dimensions of $L(\lambda)$ for the following dominant weights
@@
$\dim L(\lambda_{\ell-2}+\lambda_{\ell-1})\geq l(\lambda_{\ell-3}+\lambda_{\ell})$ $=\frac{\ell^{4}(1-2/\ell)(1-1/\ell)(1+1/\ell)}{6}>\frac{\ell^{4}(1-2/\ell)(1-1/\ell)}{6}>\ell^{4}/12>\ell^{4}/16,$
since $(1-2/\ell)(1-1/\ell)\geq((1-2/15)(1-1/15)>1/2$ for any $\ell\geq 15$. Note that $\dim L(\lambda_{\ell-2}+\lambda_{\ell-1})$ can be computed by using [13, Lemma 2.3.3], however as can be seen in our argument such a strong result is not necessary for our purposes here. In the same way, we can rule out the second case by using that $\lambda_{\ell-3}+2\lambda_{\ell}\prec\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$. Note that $\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible by appealing to Example 3.3 and $\lambda_{\ell-2}$ is admissible by Proposition 3.1.
@@
If $a_{\ell-1}\geq 2$ we note again that $\lambda_{\ell-2}$ and $\lambda_{\ell}$ appear in the writing of the dominant weight $\lambda-\alpha_{\ell-1}$, which is impossible as was shown previously. Thus, we reduce to $a_{\ell-1}=1$, i.e.,
@@
In this case we may prove that $a_{i}=0$ for $2\leq i\leq 4$. Here, many cases arise for the $\mathcal{W}$-stabilizer of $\lambda$ according to the vanish for some of these $a_{i}$ ($2\leq i\leq 4)$. For instance, if $a_{2}\neq 0$ and $a_{3}=a_{4}=0$ we see that $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{1}\times A_{\ell-4}$ or $A_{\ell-4}$ according to $a_{1}=0$ or not. Thus, we may verify that $\dim L(\lambda)>\ell^{4}/16$, i.e., $\lambda$ is not admissible. Indeed, in any case we may deduce that the dimension $L(\lambda)$ is bigger than the bound $\ell^{4}/16$. This implies
@@
Reasoning as above, one verifies that $a_{1}$ and $a_{\ell}$ cannot be simultaneously nonzero. In fact, assume first that $a_{1}\geq 2$ we obtain the dominant weight
@@
which is ruled out by computations as above. On the other hand, if $a_{\ell}\geq 2$ then $2\lambda_{\ell-1}$ appears in the writing of $\lambda-\alpha_{\ell}$, which is impossible too. Moreover, we may discard $\lambda=\lambda_{1}+\lambda_{\ell-1}+\lambda_{\ell}$ by arguing as in Example 3.3. This lefts us the dominant weights $\lambda_{1}+\lambda_{\ell-1}$, $\lambda_{\ell-1}+\lambda_{\ell}$ and $\lambda_{\ell-1}$ which are admissible by Proposition 3.1. Finally suppose that $a_{\ell-3}=a_{\ell-2}=a_{\ell-1}=0$ and $a_{\ell}\neq 0$, i.e., we will deal with
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+a_{3}\lambda_{3}+a_{4}\lambda_{4}+a_{\ell}\lambda_{\ell}.$
@@
Observe that if $a_{2}\neq 0$ then $a_{1}=0$ by using a dual reasoning to the previous case. This implies that $\lambda=\lambda_{2}+a_{\ell}\lambda_{\ell}$, which gives just one new admissible weight $\lambda=\lambda_{2}+\lambda_{\ell}$. Now, if $a_{2}=0$ then we have that $\lambda=a_{1}\lambda_{1}+a_{\ell}\lambda_{\ell}$ and we get the new admissible weights $\lambda_{1}+\lambda_{\ell}$, $2\lambda_{1}+\lambda_{\ell}$, $\lambda_{1}+2\lambda_{\ell}$ and $a_{\ell}\lambda_{\ell}$ for $a_{\ell}=1,2,3,4$. This concludes the proof concerning admissibility. Finally, Proposition 3.1, Remark 3.2, [24, Theorem 5.1] and [13, Lemma 2.3] give us the exact dimensions. ∎
@@ 4. Multiplicities and dimension in finite classical groups
In this section we shall determine all the admissible weights $\lambda$ for the rest of finite classical groups, that is, those such that $\dim L(\lambda)<\ell^{4}$. To this end, we will need to compute and bound a series of multiplicities $m_{\lambda}(\mu)$ for $\mu\preceq\lambda$.
@@ 4.1. Type $B_{\ell}$
We shall begin the desired classification of admissible weights by considering $G$ of type $B_{\ell}$. In order to compute the multiplicities involved in this section, we start by considering $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{\ell}\}$ the simple roots system (of type $B_{\ell}$) given by
@@
[…]
@@
[…]
@@
Let $p$ be an odd prime and $V$ be the standard representation of $G(\overline{\mathbb{F}}_{p})$ of dimension $2\ell+1$. For all $\ell\geq 12$, the weight $\lambda=4\lambda_{\ell}$ is admissible. To see that, it suffices to show that $\dim\operatorname{Sym}^{4}(V)$ is less or equal to $\ell^{4}$ since $L(4\lambda_{\ell})\subseteq\operatorname{Sym}^{4}(V)$. Indeed, we observe that
$\dim\operatorname{Sym}^{4}(V)=\frac{(2\ell+1)(\ell+1)(2\ell+3)(\ell+2)}{6}=\ell^{4}\cdot\frac{(2+1/\ell)(1+1/\ell)(2+3/\ell)(1+2/\ell)}{6}\\ \leq\ell^{4},$
@@
Determining all the admissible weights $\lambda$ will imply to bound several multiplicities. We start with the following lemmas.
@@
For all $\ell\geq 12$ and $p\neq 2$, a bound for the multiplicities of all dominant weights $\mu\preceq\lambda_{\ell-i}$, $i=2,3$, are given in the following table:
highest weight dominant weight $\mu\preceq\lambda$  a bound of $m_{L(\lambda)}(\mu)$
$\lambda_{\ell-2}$ $\lambda_{\ell-2}$ 11
$\lambda_{\ell-1}$ 11